OpenStudy (anonymous):

Calculate -7/4 divide(7/6)

OpenStudy (anonymous):

\[\frac{-\frac{7}{4}}{\frac{7}{6}} = -\frac{7}{4} * \frac{6}{7} = - \frac{42}{28} = - \frac{3}{2}\]